OpenStudy (anonymous):

If i have 36 = 8x^2 divide by 8 on both sides I get 4.5 = x^2 How do I get rid of the x^2?

OpenStudy (curry):

naturla log

OpenStudy (anonymous):

I dont know what that is. Can you show me please?

OpenStudy (curry):

answer would be 2.12

OpenStudy (curry):

natural log might be bit too advanced here is a nother way

OpenStudy (curry):

\[\sqrt[2]{4.5}\]

OpenStudy (anonymous):

x=√(5/2)

OpenStudy (curry):

dont for got medal

OpenStudy (anonymous):

dang, thats what i thought. now my whole problem looks even more screwy. agh!

OpenStudy (anonymous):

You need to do this: \[(4.5)^{1/2}=(x ^{2})^{1/2}\] This will cut the square amd you will have \[\sqrt[2]{4.5} = x\]

OpenStudy (zarkon):

\[36 = 8x^2\] \[\frac{36}{8}=x^2\] \[\frac{9}{2}=x^2\] \[x=\pm\sqrt{\frac{9}{2}}\]

OpenStudy (phi):

You could take the root of 9. Still not very pretty...

OpenStudy (zarkon):

one could simplify it to \[x=\pm\frac{3\sqrt{2}}{2}\]

OpenStudy (anonymous):

See now when I plug that into the original equation, im lost: 4x^2 + 9y^2 = 36 4(2.12) = 9Y^2 = 36 8.48 + 9y^2 = 36

OpenStudy (anonymous):

I feel like this cant be right?

OpenStudy (zarkon):

what did they want you to do with the original equation?

OpenStudy (anonymous):

find the dimensions of a rectangle with the greatest area that can be inscrubed in the ellipse

OpenStudy (anonymous):

i folowed all the steps, and i still want to shoot my brains out. :(

OpenStudy (zarkon):

lol...don't do that

OpenStudy (anonymous):

Im too cute for this crap, anyone want to support me for life? *sigh*... jk

OpenStudy (zarkon):

What are the din=mentions of your rectangle in terms of x,y?

OpenStudy (zarkon):

the center of the rectangle will be the center of the ellipse then let x,y be the coordinate of the intersection of the ellipse and the rectangle in the first quadrent

OpenStudy (zarkon):

then the area of the rectangle is \[A=2x\cdot 2y\]

OpenStudy (curry):

zarkon can you help me wiht my question

OpenStudy (curry):

zarkon can you help me wiht my question

OpenStudy (zarkon):

you want to maximize this subject to the constraint \[4x^2+9y^2=36\]

OpenStudy (curry):

i poseted it

OpenStudy (zarkon):

i'll look in a few mins

OpenStudy (anonymous):

ok my notes say: solve for y, so i got y=sqrt 36-4x^2 / 3

OpenStudy (zarkon):

\[y=\frac{\sqrt{36-4x^2}}{3}\] yes

OpenStudy (anonymous):

yeah, exactly that

OpenStudy (anonymous):

then i get = 1/3x * sqrt 36-4x^2 to get rid of the denominator

OpenStudy (zarkon):

?

OpenStudy (anonymous):

thats how he did it in all the problems, we do the product rule next. wrong?

OpenStudy (zarkon):

go back to \[A=2x\cdot 2y\] and replace \[y\] with what you found it to be equal to.

OpenStudy (zarkon):

\[A=2x\cdot 2\frac{\sqrt{36-4x^2}}{3}\]

OpenStudy (zarkon):

now this is a function of one variable. take the derivative ...set it to zero and solve for x

OpenStudy (anonymous):

ok im gonna try that, then im going to shame my brain for being so dumb. Thank you!!

OpenStudy (zarkon):

np ;)

OpenStudy (phi):

Hmmm, I think you got it right, especially if you use the form \[x=\frac{3\sqrt{2}}{2}\] which is exact, rather than 2.12 which is approximate

OpenStudy (phi):

Of course, you now must solve for y. and then the area of the rectangle is 4xy

OpenStudy (phi):

As a check, the area of an ellipse is \[Area= \pi a b\] which for a=3 and b=2, gives about 18.85. The rectangle must have less area than this.

OpenStudy (zarkon):

I get \[x=\frac{3\sqrt{2}}{2}\] when I maximize the area :)

OpenStudy (phi):

Looks very familiar!